Suppose a car traveling with a speed v skids a distanced after its brake locks. Calculate how far it would skid if it was traveling at a speed 2v when its brake got locked. [ Ans : 4d, Use the equation :v(=v;L 2as]
Answer #1

We will use the following equation of motion,

V2 = U2 + 2aS

where

V = final velocity ( it will be zero in our case )

U = initial velocity

a = acceleration ( will be negative in our case as car is decelerating )

so final equation

0 = U2 - 2aS

so a = U2 / 2S

now case 1 where U = v and S = d

so a = v2 / 2d

now Case 2 where U = 2v so

a = (2v)2 / 2S

from case 1 v2 / 2d = (2v)2 / 2S

so S = 4d

feel free to drop comment if you have any doubt.